# Advocating for the Silent: Enhancing Federated Generalization for Non-Participating Clients

### v5

@@ IV-A Maximum Entropy Aggregation
In this paper, we only consider the label skew scenario in FL, the empirical entropy can be thus calculated via $\hat{H}_{i}=-\sum_{y\in\mathcal{Y}}\frac{\sum_{j}\mathbb{I}_{y=y_{j}^{i}}}{n_{i}}\log\frac{\sum_{j}\mathbb{I}_{y=y_{j}^{i}}}{n_{i}}$, where $y_{j}^{i}$ is the label of $j$-th sample of local dataset $S_{i}$ of client $i$ and $\mathbb{I}$ denotes the indicator function. The proposed maximum entropy aggregation can be applied to other distribution shift scenarios if we can estimate the empirical entropy of data source [41]. How to leverage this aggregation method to benefit the federated generalization for other scenarios is out of the scope of this paper.
@@ IV-B2 Convex hull construction-based client selection
Notice that the convex hull of a point set $C$ represents the smallest convex set that encompasses all the points in $C$. The definition of the convex hull is visually depicted in Fig. 2.
@@ V Experiment
We compare the proposed empirical entropy-based weighting method with two baselines, namely FedALRC [40] and FedGAMMA [38]. Additionally, we compare two presented client selection methods with the active client selection proposed in [54]. The experimental results are depicted in Fig. 7.
@@
In Fig. 7 (a), we employ a third-order moving average to present the results. We observe that the proposed convex hull construction-based client selection method performs worse than active selection in the later stage of FL. However, the proposed selection method based on minimax gradient similarity outperforms this baseline. This difference in performance can be attributed to the sensitivity of building convex hulls for complex models such as ResNes models, while using cosine similarity to measure gradient similarity demonstrates increased robustness in this situation. Fig. 7 (b) illustrates that the proposed empirical entropy-based weighting method outperforms other baseline methods. This superiority is due to the fact that these baseline methods only consider enhancing I.I.D. generalization performance and overlook potential non-participating clients. In contrast, our presented empirical entropy-based weighting method can improve the generalization capacity of models by identifying representative data sources.
@@ VI Conclusion
This paper addresses the generalization issue in FL by exploring whether a global model trained by participating clients is capable of performing well for non-participating clients in the presence of heterogeneous data. To capture the generalization gap in FL, we propose an information-theoretic generalization framework that takes into account both the information entropy of local distribution and the discrepancy between different distributions. Leveraging this framework, we are able to identify the generalization gap and further propose an empirical entropy-based weighting aggregation method, as well as two gradient similarity-based client selection methods. These methods aim to enhance federated generalization for providing better service for non-participating clients through distribution diversification. Numerical results corroborate our theoretical findings, demonstrating that the proposed approaches can surpass baselines.

### v6

@@ IV-A Maximum Entropy Aggregation
In this paper, the empirical entropy can be thus calculated via $\hat{H}_{i}=-\sum_{y\in\mathcal{Y}}\frac{\sum_{j}\mathbb{I}_{y=y_{j}^{i}}}{n_{i}}\log\frac{\sum_{j}\mathbb{I}_{y=y_{j}^{i}}}{n_{i}}$, where $y_{j}^{i}$ is the label of $j$-th sample of local dataset $S_{i}$ of client $i$ and $\mathbb{I}$ denotes the indicator function. The proposed maximum entropy aggregation can be applied to other distribution shift scenarios if we can estimate the empirical entropy of data source [41]. How to leverage this aggregation method to benefit the federated generalization for other scenarios is out of the scope of this paper.
@@ IV-B2 Convex hull construction-based client selection
Notice that the convex hull of a point set $C$ represents the smallest convex set that encompasses all the points in $C$. The definition of the convex hull is visually depicted in Figure 2.
@@ V Experiment
(a) Empirical entropy weighting method
(b) Convex-hull selection method
(c) Minimax selection method
Figure 7: The convergence analysis on OOD test accuracy of three proposed methods combined with robust algorithms to address noisy label challenges on CIFAR-10.
Additional experiments on verifying the capability of proposed methods combined with robust algorithms to address noisy label challenges: In the following, we conduct some simple experiments to validate the effectiveness of proposed methods combined with robust algorithms to cope with noisy labels challenges. In this part, we assess the OOD performance of a CNN model on CIFAR-10. Similarly, we partition the dataset into $100$ clients using Dirichlet distribution splitting with $\alpha$ of $0.09$. We also select $40$ clients to join in FL, while the remaining ones act as non-participating clients. For the hyper-parameters of local training, we utilize the batch size of $16$, the learning rate of $0.01$, and the local epoch of $3$.
Next, we introduce the noisy label setting considered in this part. Following [43], we adopt the common Bernoulli distribution method to determine whether a local training set is affected by label noise. To inject label noise, we employ the symmetric noise method [52, 53], which manually introduces noise into datasets using a label transition matrix $T$, where $T_{i,j}:=p(\tilde{y}=j|y=i)$. In particular, we utilize an equal probability distribution for flipping true labels into other labels, thereby employing a symmetric noise method.
We integrate two prevalent Byzantine-robust algorithms into our proposed approaches to tackle the noisy label issues. Firstly, we combine two proposed client selection methods with Trimmed-mean [54], which removes outliers for each model parameter, and calculates the mean based on the remaining ones to update the global model. We combine the total variation (TV) norm penalty method [55], which adds a TV norm-based proximal regularization term in each local objective, with the proposed empirical entropy-based weighting method.
Figure 7 illustrates the effectiveness of proposed approaches in combination with robust algorithms mentioned above in addressing label noise. Notice that the baseline denoted as "Noise free" represents performing proposed methods on datasets without noisy labels. The baseline called "FedAvg" indicates conducting presented methods using datasets affected by label noise without the use of robust algorithms to mitigate such degradation. In Figure 7 (a), we observe that the proposed empirical entropy-based weighting method is significantly impacted by noisy labels. However, upon combining it with the TV-norm penalty method, the corresponding OOD test accuracy shows some improvement. In Figure 7 (b) and (c), we find that two proposed client selection methods, when combined with the Trimmed-mean method, effectively tackle with noisy labels and achieve OOD test accuracy comparable to that of the noise-free setting in the later stages of FL. These experiments have demonstrated that proposed methods can effectively integrate robust algorithms to alleviate the impact of noisy labels on FL.
@@
We compare the proposed empirical entropy-based weighting method with two baselines, namely FedALRC [40] and FedGAMMA [38]. Additionally, we compare two presented client selection methods with active selection proposed in [58]. The experimental results are depicted in Figure 8.
@@
In Figure 8 (a), we employ a third-order moving average to present the results. We observe that the proposed convex hull construction-based client selection method performs worse than active selection in the later stage of FL. However, the proposed selection method based on minimax gradient similarity outperforms this baseline. This difference in performance can be attributed to the sensitivity of building convex hulls for complex models such as ResNes models, while using cosine similarity to measure gradient similarity demonstrates increased robustness in this situation. Figure 8 (b) illustrates that the proposed empirical entropy-based weighting method outperforms other baseline methods. This superiority is due to the fact that these baseline methods only consider enhancing I.I.D. generalization performance and overlook potential non-participating clients. In contrast, our presented empirical entropy-based weighting method can improve the generalization capacity of models by identifying representative data sources.
(a) Client selection method
(b) Weighting method
Figure 9: The convergence analysis on OOD test accuracy of empirical entropy weighting method and two proposed client selection methods compared with baselines on the NICO dataset.
Additional experiments on the NICO dataset with ResNet-18: To further validate the generalizability of our proposed methods, we perform more experiments using a more realistic dataset. Specifically, we compared our approaches with previously mentioned baselines (FedALRC and FedGAMMA) using the Non-I.I.D. Image dataset with Contexts (NICO) dataset [59] with ResNet-18. The NICO dataset, which encompasses two superclasses of Animal and Vehicle, is tailored for Non-IID image classification. This dataset labels images with both primary concepts (e.g., dog) and the contexts (e.g., on grass) in which these visual concepts appear. The average number of images per context ranges from $83$ to $215$, while the average number of images per class approximates $1300$, akin to ImageNet. For this evaluation, we focus solely on the Animal superclass, which consists of $10$ classes.
We employ experimental settings akin to previous experiments conducted on CIFAR-100, setting $\alpha$ of the Dirichlet distribution to $0.09$. The experimental results are presented in Figure 9. From Figure 9 (a), we find that the proposed Minimax client selection method still outperforms other methods. However, the method based on convex hull construction demonstrates similar or even inferior performance compared to active selection. This may be attributed to the bias introduced by using cosine similarity to construct the convex hull when dealing with complex datasets like NICO. Figure 9 (b) illustrates that the proposed empirical entropy-based weighting method also surpasses other baselines due to its effective leveraging of the intrinsic category diversity within NICO, thereby enhancing the OOD performance of trained models.
@@ VI Conclusion
This paper addresses the generalization issue in FL by exploring whether a global model trained by participating clients is capable of performing well for non-participating clients in the presence of heterogeneous data. To capture the generalization gap in FL, we propose an information-theoretic generalization framework that takes into account both the information entropy of local distribution and the discrepancy between different distributions. Leveraging this framework, we are able to identify the generalization gap and further propose an empirical entropy-based weighting aggregation method, as well as two gradient similarity-based client selection methods. These methods aim to enhance federated generalization for providing better service for non-participating clients through distribution diversification. Numerical results corroborate our theoretical findings, demonstrating that the proposed approaches can surpass baselines. In the future, our goal is to develop more comprehensive theories and approaches for tackling the challenge of balancing attention between rare examples and the overall distribution. It is imperative to ensure that the emphasis on rare cases of proposed methods does not compromise performance on common cases in practical scenarios.